teve has 35 coins consisting of dimes and nickels. if the value of the coins is $3.30, then how many of each type does he have
Answer by ewatrrr(24785)

 
Hi,
Let x and (35-x) represent the number of dimes and nickels respectively
Question states***CENTS makes sense
10x + 5(35-x) = 330
Solving for x
5x = 155
x = 31, the number of dimes. 4 nickels. (35-31)
CHECKING our Answer***
10*31 + 5*4 = 310 + 20 = 330cents or $3.30